a 5.550 kg mass is tied to a rope and swung in a horizontal circle. If the velocity of the mass is 53.5 m/s and the radius of the circle is 0.650 m, what is the centripetal acceleration and centripetal acceleration and centripetal force of the mass?
1
Expert's answer
2018-01-14T07:31:07-0500
The centripetal acceleration is
a=v^2/r=(53.5)^2/0.650=4400 m/s^2 .
The centripetal force of the mass is
F=ma=m v^2/r= 5.550 (53.5)^2/0.650=24400 N=24.4 kN.